OpenStudy (anonymous):

A right triangle has an area of 24ft^2. The dimensions of the triangle are increased by a scale factor of 6. What is the area of the new triangle? Answer choices: 48ft^2 72ft^2 144ft^2 864ft^2

OpenStudy (ddcamp):

\[Ratio\ of\ lengths=a:b \rightarrow Ratio\ of\ areas=a^2:b^2\]

OpenStudy (ddcamp):

Basically, if the length increases be a factor of 6, the area will increase by a factor of 6².

OpenStudy (anonymous):

So, it would be: ? \[24ft ^{2} \times 6 = 144ft ^{2} ?\]

OpenStudy (ddcamp):

The side lengths are increased by a factor of 6, so the area would increase by a factor of 6²=36. 24ft²×36=?ft²

OpenStudy (anonymous):

\[864ft ^{2}?\]

OpenStudy (lasttccasey):

Yes that is correct @mangorox